# CAVE: Controllable Authorship Verification Explanations

### v1

@@ Appendix D Human Evaluation
Appendix E Random Errors due to Computational Limitations
We acknowledge random errors (i.e., not systematic) that could have occurred in our results owing to limitations of language models and computational resources. Our method uses a language model to read two input documents and provide an output; however, in practice, language models have a fixed maximum length of input. This means that if the two input documents together (along with the instructions) do not fit into the input, it becomes necessary to truncate them, which leads to loss of information of their latter parts. This issue is exacerbated when we move towards using in-house models on user servers (which typically have a shorter context window than API-based models like GPT-4-Turbo). In our experiments with LLaMa-3-8B, we used just 1 GPU (to have a realistic user environment), which means that while training we had to restrict the input to be of maximum length 768768768768101010note: during inference, we did not limit the size of the input, but this could pose a problem if the documents became much longer, which could have affected the training for the longer documents. Further, even for generating the training data from GPT-4-Turbo for FanFiction, the excessive size of the documents made the GPT-4-Turbo generations economically infeasible; we cut our the input to to a maximum of 300 words per document, which could have created errors/incomplete rationales that eventually propagated to our downstream training of LLaMa-3-8B.

### v2

@@ Appendix F Human Evaluation
Appendix G Random Errors due to Computational Limitations
We acknowledge random errors (i.e., not systematic) that could have occurred in our results owing to limitations of language models and computational resources. Our method uses a language model to read two input documents and provide an output; however, in practice, language models have a fixed maximum length of input due to model limitations as well as GPU memory limitations. This means that if the two input documents together (along with the instructions) do not fit into the input, it becomes necessary to truncate them, which leads to loss of information of their latter parts. This issue is exacerbated when we move towards using in-house models on user servers (which typically have a shorter context window than API-based models like GPT-4-Turbo). In our experiments with LLaMa-3-8B, we used just 1 GPU (to have a realistic user environment), which means that while training we had to restrict the input to be of maximum length 768768768768121212note: during inference, we did not limit the size of the input, but this could pose a problem if the documents became much longer, which could have affected the training for the longer documents. Further, even for generating the training data from GPT-4-Turbo for FanFiction, the excessive size of the documents made the GPT-4-Turbo generations economically infeasible; we cut our the input to to a maximum of 300 words per document, which could have created errors/incomplete rationales that eventually propagated to our downstream training of LLaMa-3-8B.
Appendix H Out-of-domain inference
Test $\rightarrow$ IMDb62 Blog-Auth FanFiction
Train $\downarrow$ Acc. Cons. Acc. Cons. Acc. Cons.
IMDb62 57.7 0.95 58.7 0.92
Blog-Auth 73.8 0.93 50.6 0.74
FanFiction 68.6 0.70 57.4 0.83
Table 14: Out of domain performance by single-dataset Cave models